OpenStudy (anonymous):

How would I solve this? |2y - 2| = 15

Parth (parthkohli):

Perhaps consider this definition:\[|x| = y \iff x = -y \ \ \text{or} \ \ x = y \rm \]

OpenStudy (anonymous):

Im sorry you lost me

Parth (parthkohli):

Read the definition again and think about what it means.

OpenStudy (anonymous):

but what do the big lines mean thats whats losing me

Parth (parthkohli):

That is the absolute value. Can you please Google it up?

OpenStudy (anonymous):

yup not a problem

Parth (parthkohli):

Thanks :)

OpenStudy (anonymous):

The actual magnitude of a numerical value okay so that mean that it would be 2y+2=15?

OpenStudy (anonymous):

@hba can you help me please and thank you

OpenStudy (hba):

Well @ParthKohli has clearly shown you :)

OpenStudy (hba):

See his first comment.

OpenStudy (hba):

Either, Or, 2y-2=15 2y-2=-15

Parth (parthkohli):

Look:\[|2y - 2| = 15 \iff 2y - 2 = -15 \ \ \text{or} \ \ 2y - 2 = 15\]

OpenStudy (hba):

I guess now you can solve :)

Parth (parthkohli):

Solve both equations individually so you get something like this:\[y = \cdots \ \ \text{or} \ \ y = \cdots\]

OpenStudy (anonymous):

so it would give me y= -6.5 or y= 8.5

Parth (parthkohli):

Yuppity yuppity yup.

OpenStudy (hba):

Exactly.

OpenStudy (anonymous):

Thank you guys so much

Parth (parthkohli):

You're welcome =)